OpenStudy (anonymous):

Simplify 3x2(7x2+ 2x - 4) 21x2+ 6x - 12 21x2+ 6x3- 12x 21x4+ 6x3- 12x2 21 + 6x - 12x2 to me it looks to be the first one. ?

OpenStudy (anonymous):

i just dont know how you would get the 6x

OpenStudy (anonymous):

look more closely at the 3rd one

OpenStudy (anonymous):

Distributively multiply 3x^2 to all the terms inside the parentheses: (3x^2 * 7x^2) + (3x^2 * 2x) + (3x^2 * -4) (21x^4) + (6x^3) + (-12x^2) 21x^2 + 6x^3 - 12x^2

OpenStudy (anonymous):

The distributive property http://en.wikipedia.org/wiki/Distributive_property is HUGE in learning algebra

OpenStudy (anonymous):

Dang, I keep having typos: 21x^4 + 6x^3 - 12x^2 is the final answer

OpenStudy (aripotta):

i was just about to ask about that lol

OpenStudy (anonymous):

Ahhh, okay i was confused. thank you.

OpenStudy (anonymous):

Both of you!

OpenStudy (anonymous):

Also it's worth stating that multiplying exponents gives the sum of the powers. x^a * x^b = x^(a+b). To illustrate it: x^2 * x^3 = (x * x) * (x * x * x) = x * x * x * x * x = x^5